OpenStudy (firejay5):

Tickets for the Prestonville High School Drama Club's spring play cost $4 for adults and $3 for students. In order to cover expenses, at least $2,000 worth of tickets must be sold. 35. Write an inequality that describes the situation. 36. Graph the inequality. 37. If 180 adults and 465 students ticket are sold, will the club cover its expenses.

OpenStudy (anonymous):

Let x be the number of adult tickets sold. Let y be the number of student tickets sold. x+ y is greater than or equal to 2000. Sorry, my equation button doesn't work for some reason.

OpenStudy (anonymous):

Graph this using table of values or algebra but I prefer table of values as follows

OpenStudy (anonymous):



OpenStudy (anonymous):

You can also use the X-intercept and Y intercept to graph which is the easiest.

OpenStudy (anonymous):

Now graph the points from the table of values to create the inequality. ( Don't forget to shade!)

OpenStudy (anonymous):

Let x be the number of adult tickets sold. Let y be the number of student tickets sold. 180 adults and 465 students. Total = 180x + 465y Total = 180(4) + 465(3) Total = 720 + 1395 Total = 2115 It is enough to cover the costs.

OpenStudy (firejay5):

 there was some numbers missing

OpenStudy (anonymous):

All ym numbers are there O_O

OpenStudy (firejay5):

does the equation look like x + y > or equal to 2,000

OpenStudy (anonymous):

Yeah.

OpenStudy (firejay5):

I understand man if it doesn't work

OpenStudy (firejay5):

How did you get 4 and 3

OpenStudy (firejay5):

@jiji501 How would you graph it and shade it

hero (hero):

You should check some youtube videos to learn how to graph

OpenStudy (firejay5):

I know how to graph it's just the equation for it

OpenStudy (firejay5):

I have 180x + 465y > or equal to 2,000

OpenStudy (firejay5):

@Hero How would the equation look for me to put it into a graphing calculator

hero (hero):

4x + 3y ≥ 2000

hero (hero):

Now, Isolate y, then graph